OpenStudy (anonymous):

The angle made by a wasp's wings with the horizontal is given by the equation \[\theta=0.4\sin600t\] where t is the time in seconds. How many times a second do its wings oscillate? Find an ezpression for \[\theta=0.4\sin600t\], the angular velocity, in radians per second. What is the value of theta when the angular velocity has (a) its greatest magnitude (b) its smallest magnitude

OpenStudy (anonymous):

sorry, find an expressions for \[d \theta \over d t\]

OpenStudy (campbell_st):

\[d \theta/dt = 0.4 \times 600 \cos 600t\] = 240cos 600t

OpenStudy (anonymous):

I understand how you got that.

OpenStudy (anonymous):

But what about the first one?

OpenStudy (campbell_st):

the oscillation is n 600...I think

OpenStudy (campbell_st):

oops the oscillation is since the period is 2pi/600

OpenStudy (anonymous):

It says 300/pi